# May the Feedback Be with You! Unlocking the Power of Feedback-Driven Deep Learning Framework Fuzzing via LLMs

### v2

@@ Abstract.
To fill this gap, we propose FUEL to effectively utilize the feedback information, which comprises two Large Language Models (LLMs): analysis LLM and generation LLM. Specifically, analysis LLM infers analysis summaries from feedback information, while the generation LLM creates tests guided by these summaries. Furthermore, based on multiple feedback guidance, we design two additional components: (i) a feedback-aware simulated annealing algorithm to select operators for test generation, enriching test diversity. (ii) a program self-repair strategy to automatically repair invalid tests, enhancing test validity. We evaluate FUEL on the two most popular DL frameworks, and experiment results show that FUEL can improve line code coverage of PyTorch and TensorFlow by 9.15% and 14.70% over state-of-the-art baselines (e.g., TitanFuzz and WhiteFox). By the time of submission, FUEL has detected 104 previously unknown bugs for PyTorch and TensorFlow, with 93 confirmed as new bugs, 49 already fixed, and 5 assigned CVE IDs.
@@ 1. Introduction
To alleviate the above issue, the success of Large Language Models (LLMs) (OpenAI, 2023a, b) in certain fields (Fan et al., 2023; Li et al., 2024) has prompted recent research to introduce LLMs into fuzzing. One of the first classical studies is TitanFuzz (Deng et al., 2023), whose key insight is that LLMs can implicitly learn both language syntax and intricate DL API constraints for DL program generation. Inspired by TitanFuzz, some follow-up studies (Deng et al., 2024b; Lyu et al., 2024; Yang et al., 2024, 2025) are proposed to explore the potential of LLMs in fuzzing. The core idea behind these studies is to leverage the strong generative abilities of LLMs to generate or mutate inputs directly. Under such a trend, the “magnificent edifice” of fuzzing has already been built, but one dark cloud still hangs in the air: Have we fully exploited the potential of feedback information during the fuzzing loop, especially the semantically rich ones? Before answering this question, we first analyze the limitations of existing studies:
@@
Limitation #2: Coarse-grained feedback analysis. Each test execution in fuzzing yields various semantic information, which is usually a mixture of programming language and natural language. state-of-the-art (SOTA) feedback-driven fuzzers still distill feedback information into coarse signals (e.g., coverage increase), thereby discarding the fine-grained, context-rich semantic information. Even PromptFuzz (Lyu et al., 2024), which mitigates coverage plateaus, only focuses on whether new code is covered, ignoring semantics from concrete code snippets. This coarse-grained manner prevents the fuzzer from reasoning about the actual runtime state of the system under test (SUT). As Figure 1 shows, coverage information pinpoint specific code snippets, bug reports contain symptom descriptions, and exception logs provide detailed stack traces. Each feedback category can be effectively processed by LLMs, which transform raw feedback into actionable insights to guide subsequent test generation.
@@
  • Approach. We implement FUEL as a practical fuzzer for DL framework testing, which utilizes feedback information and repairs invalid tests. It can be easily extended by customizing more types of feedback. FUEL is aviliable at https://github.com/NJU-iSE/FUEL.
  • Evaluation. Our experimental results show that the tests generated by FUEL achieve superior performance compared to SOTA baselines (TitanFuzz and WhiteFox), improving line coverage on two DL frameworks (PyTorch and TensorFlow) by 9.15% and 14.70%, respectively. Furthermore, ablation studies validate the contribution and effectiveness of each component in FUEL ’s design.
  • Real-World Contribution. So far, FUEL has detected 104 previously unknown bugs, with 93 already confirmed, 49 already fixed, and 5 assigned CVE IDs. 14 detected bugs were labeled as high priority, and one issue was labeled as utmost priority.
@@ 2.1. DL Framework Testing
Previous studies on DL framework testing can be classified into API-level fuzzers (Wei et al., 2022a; Xie et al., 2022; Deng et al., 2023) and model-level fuzzers (Wang et al., 2020; Liu et al., 2023a; Yang et al., 2024). API-level fuzzers target testing each DL framework API by infilling its parameters. Notably, a DL framework API usually corresponds to a DL operator. For instance, torch.nn.Conv2D, a PyTorch API (corresponding to 2-D convolution operator) can be filled with some parameters by API-level fuzzers, including kernel_size, stride, and padding. Then, a 4-D tensor will be created as input for torch.nn.Conv2D. Lastly, comparing the outputs on different hardware (e.g., CPU and GPU) to detect the potential bugs.
@@ 3.1. Prompting LLM for Test Generation
FUEL, while generally employing a feedback-driven fuzzing approach, is designed to generate tests without feedback guidance in two key scenarios: (i) In the initial iteration of fuzzing, $\mathcal{M_{G}}$ performs purely stochastic generation based on predefined prompts, as currently no prior feedback exists for guidance. (ii) Upon model repair failure: FUEL incorporates a collaborative self-repair strategy (as introduced in § 3.3 and § 3.5) that engages both $\mathcal{M_{A}}$ and $\mathcal{M_{G}}$. When this self-repair fails in the current iteration, $\mathcal{M_{G}}$ temporarily disregards previous feedback and reverts to a non-guided generation in the next iteration cycle. The motivation behind this strategy is to prevent FUEL from being trapped in an infinite repair loop (i.e., attempt successive iterations persistently but fail to resolve preceding errors).
@@ 3.2. Differential Testing and Oracles
Differential Testing. Differential testing has been proven to be an effective way to detect bugs in DL frameworks (Liu et al., 2023a). The DL model generated by LLM is passed to the DL framework for execution. We use differential testing to execute tests and subsequently identify bugs. In this work, we focus on the compilation feature of DL frameworks and design the cross-check mode on two different backends as follows.
@@
  • Compiler backend has been the most important feature of DL frameworks since version 2.0, pushing the performance of DL frameworks to new heights. As an emerging technology, the compiler backend is still in rapid development, and a larger number of bugs have not been detected (Yang et al., 2024). PyTorch Inductor and TensorFlow XLA are compiler backends for PyTorch and TensorFlow, respectively.
@@
Execution and cross-checking on the two backends can be classified into three distinct states (Figure 1): (i) Cross-checking results violate the oracles, and we consider this a potential bug in DL frameworks and collect bug reports. (ii) If a DL model crashes on both backends, then we consider this an invalid model and collect exception logs. (iii) Only when execution results of both backends do not violate oracles or crash, do we consider this a bug-free and valid model, and collect code coverage from it.
@@ 3.3. Prompting LLM for Analysis Summary
To address the above challenges, we introduce a $\mathcal{M_{A}}$ to provide clear, explicit, and concise analysis summary for $\mathcal{M_{G}}$. $\mathcal{M_{A}}$ infers explanation, reasons, and next testing strategy from feedback step by step. Figure 6 shows the analysis prompt template of FUEL, which begins with a contextual description: “Shown here is a [LIBRARY] model along with input tensors from the last iteration.” The core components consist of [MODEL CODE], where [MODEL CODE] represents a DL model (introduced in § 2.1). Next, $\mathcal{M_{A}}$ will select different templates based on different types of feedback, and the feedback information will be subsequently filled into the corresponding placeholders (i.e., [COVERAGE], [BUG], and [EXCEPTION]). Immediately following, the instruction makes the $\mathcal{M_{A}}$ analyze the feedback information semantically. Lastly, $\mathcal{M_{A}}$ infers explanation, reasons, and next testing strategy in [PLACEHOLDER...]. This progressive step provides explicit guidance for test generation in subsequent iterations. More specific examples of analysis prompts will be discussed in § 5.3.
@@ 3.4. Heuristic Search for Operators Selection
While LLM can benefit from sufficient analysis summaries, it tends to use some common operators (e.g., Conv2d, ReLU, and Linear) for test generation (Deng et al., 2024b), significantly damaging model diversity (discussed in Limitation #4). This limitation motivates the implementation of RAG (Gao et al., 2023) for $\mathcal{M_{G}}$, a logical and natural enhancement that prioritizes less prevalent operators in test generation.
@@
Unfortunately, operator selection is a NP-hard problem, stemming from the combinatorial explosion inherent in DL model construction: While DL frameworks typically offer a finite operator set (numbering in the thousands), the unbounded combinations of these operators create an intractably large retrieval space. Notably, we identify that feedback signals can serve as indicators of intrinsic value, enabling the iterative optimization of operator selection through value-guided search. The idea behind employing heuristic search lies in reformulating the RAG task as a combinatorial optimization problem within a constrained space: At each iteration, we aim to select an optimal subset of operators from the finite collection that maximizes a predefined value function. This value function, dynamically determined by accumulated feedback signals, quantifies the strategic importance of operators for subsequent iterative generation.
@@
Inspired by the above observations, we adopt simulated annealing (Van Laarhoven et al., 1987) for operator selection in FUEL, a balance between overhead and performance. High-performance search algorithms, such as Monte Carlo Tree Search, incur prohibitive temporal overhead during incremental tree construction in iterative fuzzing cycles. Conversely, low-overhead search algorithms (e.g., random sampling) fail to identify high-value operators, compromising search performance. Simulated annealing emerges as an offline planning algorithm that efficiently approximates global optima (regulated through cooling rate $\gamma$) within constrained time budgets. For operator valuation, we design three metrics of operators: (i) Number of uses; (ii) Exception-triggering count; (iii) Number of new triggered line coverage. The valuation function follows an inverse relationship with the first two factors (i.e., frequency and exceptions), reflecting our emphasis on exploring underused operators. Conversely, it maintains direct proportionality to new code coverage, prioritizing operators that can trigger new code coverage. We define the final valuation function as follows: $V_{\alpha,\beta}(x,y,z)=\frac{\alpha}{\alpha+x}+\frac{\alpha}{e^{y}}+\beta-e^{-\frac{z}{100}}$, where $x$, $y$, and $z$ denote for number of use, exception-triggering count, and number of new triggered line coverage, respectively. The hyper-parameters $\alpha$ and $\beta$ are both initialized at $0.5$. This configuration ensures algorithmic stability before the fuzzing campaign, where $x$, $y$, and $z$ remain zero-initialized. Under this initialization setting, the value for all operators reduces to: $V_{\alpha,\beta}(0,0,0)=\alpha+\beta$. Substituting the predefined hyperparameters yields a uniform initial value $V_{0.5,0.5}(0,0,0)=1.0$, ensuring fair operator prioritization at the beginning of fuzzing.
@@ 4. Implementation
Operator collection from API document. As discussed in § 3, an operator set is needed for FASA. In this work, we collect core operators from two DL frameworks. Specifically, we implement an HTML crawler to collect operators from PyTorch with API prefixes torch, torch.nn, torch.nn.functional as they are core operators (PyTorch, 2021) in PyTorch 2.x. For TensorFlow, we collect operators that XLA supports from its document (TensorFlow, 2025) as we focus on the compilation feature of the DL framework in this work. Although we currently only consider the most important subset of all operators, we can easily modify the operator set to suit specific test scenarios (e.g., considering gradient-related operators to test the gradient feature of DL frameworks).
@@ 5.1. Experimental Setup
Baselines. In this work, we select two state-of-the-art (SOTA) fuzzers for DL framework testing (i.e., TitanFuzz (Deng et al., 2023) and WhiteFox (Yang et al., 2024)). Notably, we do not consider traditional fuzzers because WhiteFox has demonstrated that LLM-based fuzzers can outperform the SOTA traditional fuzzer (e.g., NNSmith (Liu et al., 2023a)) on both code coverage and bug detection ability.
@@ 5.2.1. Default fuzzing configuration
Figure 8 demonstrates the performance of FUEL against two SOTA baselines (TitanFuzz and WhiteFox), across both on PyTorch and TensorFlow. Over 1,000 iterations, FUEL achieves superior code coverage compared to both baselines, outperforming the second-best baseline TitanFuzz by 9.15% and 14.70% over PyTorch and TensorFlow, respectively. Notably, this enhanced code coverage is obtained despite FUEL employing fewer operators, indicating its effectiveness in combining different operators. This performance advantage may benefit from the model-level fuzzing paradigm of FUEL that comprehensively explores diverse graph patterns compared to API-level fuzzers. In contrast, WhiteFox exhibits relatively lower coverage, potentially attributed to its optimization-triggering pattern specialization. While it is effective for testing specific compiler optimizations, this feature inadvertently restricts model diversity through excessive template matching. Experimental results support our hypothesis that utilizing coverage information to guide test generation proves more effective than the API-level fuzzing technique or optimization-targeted approach.
@@ 5.3. Bug Study
Table 3 presents the bug detection results of FUEL over PyTorch and TensorFlow. We first analyze some representative detected bugs and then analyze the advantage of FUEL to trigger similar new bugs. Lastly, we explain why some detected bugs won’t be fixed by DL framework developers.
@@ 5.3.1. Analysis of bug examples
Figure 9(b) demonstrates a behavior inconsistency bug of TensorFlow, which is a security vulnerability detected by FUEL. The bug-inducing model consists of tf.reshape and tf.keras.layers.Conv2D. When setting padding='valid' for Conv2D, eager backend executes the model successfully while compiler backend throws the negative dimension size error. This bug can be leveraged as a security vulnerability by hackers who can require the Conv2D parameter in padding='valid' format, resulting in DoS attacks.
@@
Among the 104 bugs we have reported, five of them won’t be fixed by developers. Developers dismiss all of them because the numerical inconsistency is too small. They suggested increasing the tolerance threshold of accuracy. Subsequently, we did it in our next bug reports.

### v3

@@ Abstract.
To fill this gap, we propose FUEL to effectively utilize the feedback information, which comprises two Large Language Models (LLMs): analysis LLM and generation LLM. Specifically, analysis LLM infers analysis summaries from feedback information, while the generation LLM creates tests guided by these summaries. Furthermore, based on multiple feedback guidance, we design two additional components: (i) a feedback-aware simulated annealing algorithm to select operators for test generation, enriching test diversity. (ii) a program self-repair strategy to automatically repair invalid tests, enhancing test validity. We evaluate FUEL on the two most popular DL frameworks, and experiment results show that FUEL can improve line code coverage of PyTorch and TensorFlow by 9.15% and 14.70% over state-of-the-art baselines (e.g., TitanFuzz and WhiteFox). By the time of submission, FUEL has detected 104 previously unknown bugs for PyTorch and TensorFlow, with 93 confirmed as new bugs, 49 already fixed, and 14 assigned CVE IDs.
@@ 1. Introduction
To alleviate the above issue, recent research has introduced Large Language Models (LLMs) (OpenAI, 2023a, b) into fuzzing, motivated by their success in other fields (Fan et al., 2023; Li et al., 2024). One of the first classical studies is TitanFuzz (Deng et al., 2023), whose key insight is that LLMs can implicitly learn both language syntax and intricate DL API constraints for DL program generation. Inspired by TitanFuzz, some follow-up studies (Deng et al., 2024b; Lyu et al., 2024; Yang et al., 2024, 2025) are proposed to explore the potential of LLMs in fuzzing. The core idea behind these studies is to leverage the strong generative abilities of LLMs to generate or mutate inputs directly. Under such a trend, the “magnificent edifice” of fuzzing has already been built, but one dark cloud still hangs in the air: Have we fully exploited the potential of feedback information during the fuzzing loop, especially the semantically rich ones? Before answering this question, we first analyze the limitations of existing studies:
@@
Limitation #2: Coarse-grained feedback information analysis. Each test execution in fuzzing yields various semantic information, which is usually a mixture of programming language and natural language. state-of-the-art (SOTA) feedback-driven fuzzers still distill feedback information into coarse signals (e.g., coverage increase), thereby discarding the fine-grained, context-rich semantic information. Even PromptFuzz (Lyu et al., 2024), which mitigates coverage plateaus, only focuses on whether new code is covered, ignoring semantics from concrete code snippets. This coarse-grained manner prevents the fuzzer from reasoning about the actual runtime state of the system under test (SUT). As Figure 1 shows, coverage information pinpoints specific code snippets, bug reports contain symptom descriptions, and exception logs provide detailed stack traces. Each feedback category can be effectively processed by LLMs, which transform raw feedback into actionable insights to guide subsequent test generation.
@@
  • Approach. We implement FUEL as a practical fuzzer for DL framework testing, which utilizes feedback information and repairs invalid tests. FUEL is available at https://github.com/NJU-iSE/FUEL.
  • Evaluation. Our experimental results show that the tests generated by FUEL achieve superior performance compared to SOTA baselines (TitanFuzz and WhiteFox), improving line coverage on two DL frameworks (PyTorch and TensorFlow) by 9.15% and 14.70%, respectively. Additionally, ablation studies validate the contribution and effectiveness of each component in FUEL ’s design.
  • Real-World Contribution. So far, FUEL has detected 104 previously unknown bugs, with 93 already confirmed, 49 already fixed, and 14 assigned CVE IDs. 14 detected bugs were labeled as high priority, and one issue was labeled as utmost priority.
@@ 2.1. DL Framework Testing
Previous studies on DL framework testing can be classified into API-level fuzzers (Wei et al., 2022a; Xie et al., 2022; Deng et al., 2023) and model-level fuzzers (Wang et al., 2020; Liu et al., 2023a; Yang et al., 2024). API-level fuzzers target testing each DL framework API by infilling its parameters. Notably, a DL framework API usually corresponds to a DL operator. For instance, torch.Conv2D, a PyTorch API (corresponding to 2-D convolution operator) can be filled with some parameters by API-level fuzzers, including kernel_size, stride, and padding. Then, a 4-D tensor will be created as input for torch.Conv2D. Lastly, comparing the outputs on different hardware (e.g., CPU and GPU) to detect the potential bugs.
@@ 3.1. Prompting LLM for Test Generation
While generally employing a feedback-driven fuzzing approach, FUEL is designed to generate tests without feedback guidance in two key scenarios: (i) In the initial iteration of fuzzing: $\mathcal{M_{G}}$ performs purely stochastic generation based on predefined prompts, as currently no prior feedback exists for guidance. (ii) Upon model repair failure: FUEL incorporates a collaborative self-repair strategy (as introduced in § 3.3 and § 3.5) that engages both $\mathcal{M_{A}}$ and $\mathcal{M_{G}}$. When this self-repair fails in the current iteration, $\mathcal{M_{G}}$ temporarily disregards previous feedback and reverts to a non-guided generation in the next iteration cycle. The motivation behind this strategy is to prevent FUEL from being trapped in an infinite repair loop (i.e., attempt successive iterations persistently but fail to resolve preceding errors).
@@ 3.2. Differential Testing and Oracles
Differential Testing. Differential testing has been proven to be an effective way to detect bugs in DL frameworks (Liu et al., 2023a). The DL model generated by LLM is executed on different backends to identify bugs. In this work, we focus on the compilation feature of DL frameworks and design the cross-check mode on two different backends as follows.
@@
  • Compiler backend has been the most important feature of DL frameworks since the version 2.x era, pushing the performance of DL frameworks to new heights. As an emerging technology, the compiler backend is still in rapid development, and has become an important backend to compile DL models (especially LLMs) (Yang et al., 2024). PyTorch Inductor and TensorFlow XLA are compiler backends for PyTorch and TensorFlow, respectively.
@@
Execution and cross-checking on the two backends can be classified into three distinct states (Figure 1): (i) Cross-checking results violate the oracles, indicating a potential bug in DL frameworks, then we collect bug reports. (ii) If a DL model crashes on both backends, it is an invalid model and we collect exception logs. (iii) Only when execution results of both backends do not violate oracles or crash, it is a bug-free and valid model, and code coverage is collected.
@@ 3.3. Prompting LLM for Analysis Summary
To address the above challenges, we introduce a $\mathcal{M_{A}}$ to provide clear, explicit, and concise analysis summary for $\mathcal{M_{G}}$. $\mathcal{M_{A}}$ infers explanation, reasons, and next testing strategy from feedback step by step. Figure 6 shows the analysis prompt template of FUEL, which begins with a contextual description: “Shown here is a [LIBRARY] model along with input tensors from the last iteration.” The core components consist of [MODEL CODE], which represents a DL model (introduced in § 2.1). Next, $\mathcal{M_{A}}$ selects different templates based on different types of feedback, and the feedback information will be subsequently filled into the corresponding placeholders (i.e., [COVERAGE], [BUG], and [EXCEPTION]). Immediately following, the instruction makes the $\mathcal{M_{A}}$ analyze the feedback information semantically. Finally, $\mathcal{M_{A}}$ infers explanation, reasons, and next testing strategy in [PLACEHOLDER...]. This progressive step provides explicit guidance for test generation in subsequent iterations. More specific examples of analysis prompts will be discussed in § 5.3.
@@ 3.4. Heuristic Search for Operators Selection
While LLM-based approaches can benefit from sufficient analysis summaries, they tend to use some common operators (e.g., Conv2d, ReLU, and Linear) for test generation (Deng et al., 2024b), which significantly limits model diversity (discussed in Limitation #4). This limitation introduces the implementation of RAG (Gao et al., 2023) for $\mathcal{M_{G}}$, a logical and natural enhancement that prioritizes less prevalent operators in test generation.
@@
Unfortunately, operator selection is a challenging NP-hard problem, stemming from the combinatorial explosion inherent in DL model construction: While DL frameworks typically offer a finite operator set (numbering in the thousands), the unbounded combinations of these operators create an intractably large retrieval space. Notably, we identify that feedback signals can serve as indicators of intrinsic value, enabling the iterative optimization of operator selection through value-guided search. The idea behind employing heuristic search lies in reformulating the RAG task as a combinatorial optimization problem within a constrained space: At each iteration, the algorithm selects an optimal subset of operators from the finite collection that maximizes a predefined value function. This value function, dynamically determined by accumulated feedback signals, quantifies the strategic importance of operators for subsequent iterative generation.
@@
Inspired by the above idea, we adopt simulated annealing (Van Laarhoven et al., 1987) for operator selection in FUEL, a balance between overhead and performance. High-performance search algorithms, such as Monte Carlo Tree Search, incur prohibitive temporal overhead during incremental tree construction in iterative fuzzing cycles. Conversely, low-overhead search algorithms (e.g., random sampling) fail to identify high-value operators, compromising search performance. Simulated annealing emerges as an offline planning algorithm that efficiently approximates global optima (regulated through cooling rate $\gamma$) within constrained time budgets. For operator valuation, we design three metrics of operators: (i) Number of uses; (ii) Exception-triggering count; (iii) Number of new triggered line coverage. The valuation function follows an inverse relationship with the first two factors (i.e., frequency and exceptions), reflecting our emphasis on exploring underused operators. Conversely, it maintains direct proportionality to new code coverage, prioritizing operators that can trigger new code coverage. We define the final valuation function as follows: $V_{\alpha,\beta}(x,y,z)=\frac{\alpha}{\alpha+x}+\frac{\alpha}{e^{y}}+\beta-e^{-\frac{z}{100}}$, where $x$, $y$, and $z$ denote for number of use, exception-triggering count, and number of new triggered line coverage, respectively. The hyper-parameters $\alpha$ and $\beta$ are both initialized at $0.5$. This configuration ensures algorithmic stability before the fuzzing campaign, where $x$, $y$, and $z$ remain zero-initialized. Under this initialization setting, the value for all operators reduces to: $V_{\alpha,\beta}(0,0,0)=\alpha+\beta$. Substituting the predefined hyperparameters yields a uniform initial value $V_{0.5,0.5}(0,0,0)=1.0$, ensuring fair operator prioritization at the beginning of fuzzing.
@@ 4. Implementation
Operator collection from API document. As discussed in § 3, an operator set is needed for FASA. In this work, we collect core operators from two DL frameworks. Specifically, we implement an HTML crawler to collect operators from PyTorch with API prefixes torch, torch.nn, and torch.nn.functional as they are core operators (PyTorch, 2021) in PyTorch 2.x. For TensorFlow, we collect operators that XLA supports from its document (TensorFlow, 2025) as we focus on the compilation feature of the DL framework in this work. Although we currently only consider the most important subset of all operators, we can easily modify the operator set to suit specific test scenarios (e.g., considering gradient-related operators to test the gradient feature of DL frameworks).
@@ 5.1. Experimental Setup
Baselines. In this work, we select two SOTA fuzzers for DL framework testing (i.e., TitanFuzz (Deng et al., 2023) and WhiteFox (Yang et al., 2024)). Notably, traditional fuzzers are not included because WhiteFox has demonstrated that LLM-based fuzzers can outperform the SOTA traditional fuzzer (e.g., NNSmith (Liu et al., 2023a)) on both code coverage and bug detection ability.
@@ 5.2.1. Default fuzzing configuration
Figure 8 demonstrates the performance of FUEL against two SOTA baselines (TitanFuzz and WhiteFox), across both on PyTorch and TensorFlow. Over 1,000 iterations, FUEL achieves superior code coverage compared to both baselines, outperforming the second-best baseline TitanFuzz by 9.15% and 14.70% on PyTorch and TensorFlow, respectively. Notably, this enhanced code coverage is obtained despite FUEL employing fewer operators, indicating its effectiveness in combining different operators. This performance advantage may benefit from the model-level fuzzing paradigm of FUEL that comprehensively explores diverse model structures compared to API-level fuzzers. In contrast, WhiteFox exhibits relatively lower coverage, potentially attributed to its optimization-triggering pattern specialization. While it is effective for testing specific compiler optimizations, this feature inadvertently restricts model diversity through excessive template matching. Experimental results support our hypothesis that utilizing coverage information to guide test generation proves more effective than the API-level fuzzing technique or optimization-targeted approach.
@@ 5.3. Bug Study
Table 3 presents the bug detection results of FUEL over the last seven months. We first analyze some representative detected bugs and then analyze the advantage of FUEL to trigger similar new bugs. Lastly, we explain why some detected bugs won’t be fixed by DL framework developers.
@@ 5.3.1. Analysis of bug examples
Figure 9(b) demonstrates a behavior inconsistency bug of TensorFlow, which is a security vulnerability detected by FUEL. The bug-inducing model consists of tf.reshape and tf.keras.layers.Conv2D. When setting padding='valid' for Conv2D, eager backend executes the model successfully, while compiler backend throws the negative dimension size error. This bug can be leveraged as a vulnerability by requiring the Conv2D parameter in padding='valid' format, resulting in DoS attacks, which was assigned CVE-2025-55559.
@@
Among the 104 bugs we have reported, five of them won’t be fixed by developers. Developers dismiss them because the numerical inconsistency is too small. They suggested increasing the tolerance threshold of accuracy. Subsequently, we did it in our next bug reports.